OpenStudy (anonymous):

-5x+2y=6 -15x+6y=0 Im trying to find the slope of these lines can anyone assist me?

OpenStudy (anonymous):

If you turn them into the form y = mx +b it becomes easier to find the slope -5x+2y = 6 2y = 5x + 6 y = 5/2x +3 therefore, the slope is 5/2. -15x + 6y = 0 6y = 15x y = 15/6x or y = 5/2x The slope is 5/2

OpenStudy (anonymous):

@ssiddiq101 thank you!

OpenStudy (anonymous):

no worries :)

OpenStudy (anonymous):

@ssiddiq101 how do I solve by graphing x-3y=-3 & x+3y=9

OpenStudy (anonymous):

just turn them into y = mx + b form, you should be able to get the slope and y-intercept and then graph some points.

OpenStudy (anonymous):

alright